# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ Abstract
In hypothesis testing with quantum states, given a black box containing one of the two possible states, measurement is performed to detect in favor of one of the hypotheses. In postselected hypothesis testing, a third outcome is added, corresponding to not selecting any of the hypotheses. In postselected scenario, minimum error one-shot symmetric hypothesis testing is characterized in literature conditioned on the fact that one of the selected outcomes occur. We proceed further in this direction to give the set of all possible measurements that lead to the minimum error. We have given an arbitrary error-minimizing measurement in a parametric form. Note that not selecting any of the hypotheses decimates the quality of testing. We further give an example to show that these measurements vary in quality. There is a need to discuss the quality of postselected hypothesis testing. We then characterize the quality of postselected hypothesis testing by defining a new metric acceptance and give expression of acceptance for an arbitrary error-minimizing measurement in terms of some parameters of the measurement. On the set of measurements that achieve minimum error, we have maximized the acceptance, and given an example which achieves that, thus giving an example of the best possible measurement in terms of acceptance.
@@ I Introduction
Quantum hypothesis testing deals with discriminating between the hypotheses corresponding to quantum properties of the nature with plethora of applications in quantum information science [2, 3, 4]. The quantum state discrimination is an example of quantum hypothesis testing which aims to make a decision based on a measurement of a given quantum object, which can be in one of the two possible states corresponding to the null and alternative hypotheses. If the decision has to be made on the basis of measurement of a single copy of the quantum object, it is called one-shot hypothesis testing and if infinitely many copies are available, it called asymptotic hypothesis testing. When the test falsely concludes the alternative hypothesis, it is called type-1 error; when the test falsely decides in the favor of the null hypothesis, it is called type-2 error. Ideally the goal is to select a measurement and a decision rule to have both type-1 and type-2 errors arbitrarily small. However, simultaneous minimization of both error is difficult. Hence, a trade-off is usually selected. For example, in asymmetric hypothesis testing, the goal is to minimize type-2 error under some constraint on the type-1 error. Another example is the symmetric hypothesis testing where the objective is to minimize the average probability of error.
@@
In the one-shot discrimination of quantum states, the minimum possible error can be easily obtained using semi-definite programming for both symmetric and asymmetric hypothesis testing problems. In symmetric hypothesis testing, the error is characterized in a closed-form expression by the Helstrom-Holevo theorem [5, 6]. Error exponent in quantum symmetric hypothesis testing was characterized in [7, 8]. Quantum relative entropy[9] was shown to be the error exponent of asymmetric hypothesis testing in [10, 11]. Optimal measurements to achieve this error are given in [12]. In conventional quantum hypothesis testing, as described above, one of the two hypotheses is selected that concludes the presence of one of the hypotheses. Here, the key problem remains that, non-orthogonal states can not be perfectly distinguished. To avoid this problem approaches utilizing inconclusive decision rules were proposed in [13, 14, 15] which were later generalized in [16, 17, 18, 19, 20, 21, 22] focusing on maximizing confidence in the outcome of the measurement. A comprehensive survey of various quantum state discrimination strategies is given in [23, 24].
While it may not be possible to perfectly distinguish non-orthogonal states, it is possible to design measurements that include an extra outcome addition to the two outcomes corresponding to each state [25], which corresponds to not selecting any of the hypotheses (see Fig. 1). Here, the decision is only made when third outcome does not occur, hence it is termed postselected hypothesis testing. Postselected probabilities are defined as probability of any event conditioned that one of the selected outcomes is observed. The postselected error probabilities have been characterized in [25] for symmetric and asymmetric hypothesis testing, both in one-shot or asymptotic cases and derived the minimum possible error. It is also generalized to hypothesis testing for quantum channels and classical probability distributions and an example is given to show that minimum error bound is achievable in all the mentioned cases [25]. In our preliminary work [1], we define a metric acceptance signifying the probability that one of two outcomes occur and a decision is made to select one of the two hypothesis.
(a) Conventional quantum state discrimination: The measurement gives two outcomes, each corresponding to one of the possible states of the object in unknown state.
(b) Postselected quantum state discrimination: The measurement gives an addition outcome (given in red), corresponding to not selecting either of the states.
Figure 1: The unknown-state quantum object is in the state $\nu\in\{\rho,\sigma\}$. A measurement is performed to detect the state.
Motivation and Contributions: It is reasonable to wonder if finding an error minimization-measurement is sufficient in postselected hypothesis testing. Let us consider a case where there are two measurements having the same error but different probability of rejection. Essentially, even though error is same, the one having lower rejection is better. Further, while it may be possible that a measurement that minimizes the error has a high probability of having third outcome (i.e. rejection). As such measurements may practically be unusable, this brings a question if and how we can find the measurements that minimize the probability of rejection as well as the postselected error. The key contributions of present work is as follows:
@@
    We begin with deriving the condition on the measurements for it to be an error-minimizing measurement. We obtain three different conditions in three different cases depending on the prior probability of unknown state being $\rho$ (and $\sigma$). Based on the conditions, we observe that the error-minimizing measurement never makes decision in favor $\rho$ (or $\sigma$), if prior probability of unknown state being in the state $\rho$ i.e. $p_{\rho}$ is smaller (or greater) than a threshold. If it is equal to the the threshold, the error-minimizing measurement may make a decision in favor of either/both of the states. A classical example (see Example 1) is given to illustrate finding measurements from this and further showing that acceptance varies over the set of all error minimizing measurement.
@@
    We have given the set of all error-minimizing measurements in parameterized form for all the three cases. This a generalization from the literature in the sense that achievability was shown in literature but exhausting set of all such measurements is not known. We give a method to construct an arbitrary error-minimizing measurement in terms of certain parameters. The method is summarized in a table for the three cases. A quantum example in two parts (see Example 2(a) and 2(b)) is given to illustrate finding error minimizing measurements.
@@
    For an arbitrary measurement from the set of error minimizing measurements, we have derived the a closed-form expressions of acceptance. We have given a closed-form expression of maximum acceptance obtained from maximizing over set of all error-minimizing measurements in many cases. In some cases, we have given it as an optimization problem.
@@
    Then we proceed to generalize the results to the case when the support are not equal i.e., $\Pi_{\rho}\neq\Pi_{\sigma}$. In this case, we observe that the minimum-error vanishes. Then, we find the condition on an arbitrary measurement to ensure that the error vanishes. We observe that the error may vanish in three different ways and then write the set of error-minimizing measurements as union of the sets satisfying the three conditions. We discuss the property of the measurements from the three sets. We then shown that two of the sets remain empty, when $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$.
@@
    We give parameterization for an arbitrary measurements for the three sets. We then find expression of acceptance. In the end, we obtain the expression of maximum acceptance in closed form expression by taking maximum of the maximum acceptance obtained on the three sets.
The notations needed across the paper is given below in this section before describing the system model in the next section.

### v2

@@ Abstract
Quantum state discrimination refers to the determination of the unknown state of a quantum object that can be in one of two possible states, by performing a certain measurement that declares its state to be one of two states. Hence, it can be expressed as a hypothesis testing problem with null and alternative hypotheses representing the two states. Postselected state discrimination (PSD) considers measurements with an additional outcome, corresponding to not making any determination, thus rejecting the test. Thus, allowing the test not to make a decision in unreliable cases can help PSD reduce the effective error in state determination among cases when a decision is made. However, this comes at the expense of reduced probability of making a decision which we term acceptance. In this work, we focus on deriving optimal measurements that maximize acceptance while minimizing the postselected error. We first give the set of all measurements that can achieve the minimum error, along with a method to parameterize these error-minimizing measurements. We further provide an example to demonstrate that these measurements vary in acceptance and thus establish the need to choose a measurement that minimizes error and maximizes acceptance simultaneously. We derive the expression of acceptance for an arbitrary error-minimizing measurement in terms of the parameters used for parameterization proposed earlier. We find the maximum value of acceptance achievable over the set of all error-minimizing measurements. We also provide an example measurement that achieves the maximum acceptance.
@@ I Introduction
Quantum hypothesis testing involves discriminating between hypotheses corresponding to the quantum properties of nature, with numerous applications in quantum information science [40, 20, 41]. A specific instance of quantum hypothesis testing is quantum state discrimination, which aims to determine the state of a given quantum object that can exist in one of two possible states ($\rho$ and $\sigma$), based on a measurement on the object. These states can be regarded as the null and the alternative hypotheses. If the decision has to be made based on the measurement of a single copy of the quantum object, it is referred to as one-shot hypothesis testing. If infinitely many copies are available, it is called asymptotic hypothesis testing. When the test falsely concludes the alternative hypothesis, it is called type-1 error; when the test falsely decides in favor of the null hypothesis, it is called type-2 error. Ideally, the goal is to select a measurement and a decision rule with both type-1 and type-2 errors arbitrarily small. Since simultaneous minimization of the two errors is difficult, a trade-off is usually adopted. For example, in asymmetric hypothesis testing, the objective is to minimize the type-2 error under a constraint on the type-1 error. On the other hand, symmetric hypothesis testing aims to minimize the average error probability.
@@
In one-shot discrimination of quantum states, the minimum error achievable in an asymmetric hypothesis testing problem can be obtained using semi-definite programming. In symmetric hypothesis testing, the error can be expressed in a closed-form expression utilizing the Helstrom-Holevo theorem [21, 24]. In the asymptotic case, the minimum error in quantum hypothesis testing vanishes exponentially with the number of copies of states, and the error exponent was characterized in [29, 1] for the symmetric case and in [23, 30] for the asymmetric case. In particular, quantum relative entropy[39] was shown to be the error exponent of asymmetric hypothesis testing. Further, [19] presents the optimal measurements that can achieve this error. In conventional quantum hypothesis testing, as described above, one of the two hypotheses is selected. The key problem remains that non-orthogonal states cannot be distinguished with zero error via conventional quantum hypothesis testing.
To distinguish the non-orthogonal states with zero-error, prior work [14, 9, 13, 37, 4, 16, 42, 43, 7, 26, 28, 18, 17, 25, 12, 31, 8, 36, 10, 22, 3, 44] have proposed to include an extra outcome in the quantum hypothesis testing where no state is declared, effectively, rejecting both hypotheses, or equivalently declaring a failure of the test. Such discrimination strategies are termed unambiguous state discrimination (USD) (see [9, 13, 37, 4, 16, 43, 26]). For an example, consider a pair of states $\{\rho,\sigma\}$. USD aims to design a measurement such that only possible outcomes are (i) declare $\rho$, and (ii) reject the test, when the state is $\rho$ and; the possible outcomes are (i) declare $\sigma$ and (ii) reject the test when the state is $\sigma$. This is illustrated in Fig. 1 (b). In other words, in a USD, the measurement either detects the correct state or declare a decision failure, resulting in a zero error among accepted (selected) results. However, it is important to note that there is a certain probability that a decision is not made, which can be termed failure or test-rejection probability. Therefore, it is desirable to have a low failure probability while ensuring zero error in USD. While desirable, it is noteworthy that USD is not always possible, as its feasibility depends on the state pair $\{\rho,\sigma\}$ we aim to distinguish. The prior work in USD has focused on determining the condition on the state pair for USD to be feasible, determining the USD measurement under the above condition, and determining USD measurements with the minimum probability of rejection, along with derivation of specific bounds on the rejection probability [9, 13, 37, 4, 16, 26]. In Table I, we summarize related papers from the past literature along with their system model and relevant contributions. Readers are suggested to refer to [5], and [2] for a comprehensive survey of various quantum state discrimination strategies.
Ref Limited to USD? Type of states System model with key relevant contributions
[9] YES $n$ Pure states It is shown that USD is possible only if the states are linearly independent. A general form of USD measurements is obtained. A discussion on maximizing the success probability (expected acceptance) is included and illustrated by an example.
[13] YES $n$ Mixed states A lower bound on the probability of rejection (=1=1-expected acceptance) is obtained in terms of the fidelity of each possible pair of states.
[37] YES 22 Mixed states Rank-2 states in Hilbert’s space of dimension $4$ are considered for discrimination. A measurement that maximizes the success probability (expected acceptance) is numerically obtained.
[4] YES $n$ Pure states Prior probability of states is known. An upper bound is obtained on the maximum success probability (expected acceptance) and shown to be tight and achievable in some cases.
[16] YES $n$ Pure states Each pair of states has an equal and real inner product. For states with equal prior probability, for USD measurements, the success probability (expected acceptance) is maximized.
[42] UCD $n$ Pure states The paper introduces unambiguous correct detection (UCD), a different generalization of USD. A new metric (=11-postselected symmetric error) is defined and it is shown that UCD improves the metric with help of examples.
[43] YES 22 Pure states Unknown quantum states with multiple available copies are taken for discrimination, and a method to design a programmable circuit for USD is presented.
[7] More General 22 Pure states In this work, given a quantum object in one of the two possible quantum states, a new pair of quantum states with more separation (in terms of the inner product) is obtained, with partial success. The paper states the maximum probability of success (expected acceptance) for such operations and also provides a method to design such an operation.
[26] YES $n$ Pure states For pure states with known prior probabilities $\{(|\psi_{i}\rangle,\eta_{i}):i\in[1,n]\}$, USD measurements are obtained as $\Pi_{i}=p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, where reciprocal states $|\psi^{\perp}_{k}\rangle:\langle\psi^{\perp}_{k}|\psi_{i}\rangle=\delta_{ik}$. The probability of inconclusive outcomes $\left(=1-\sum_{k}\eta_{k}p_{k}\right)$ is minimized. An exact solution is obtained for $n=2,3$, and a numerical method is given for more states.
[28] YES 2 Mixed states Measurement for which error vanishes (i.e., USD measurements), and success probability (i.e., expected acceptance) is maximum, is obtained numerically by solving a set of equations, taken from a prior work.
[18] More general 2 Pure states $|\psi_{1}\rangle,|\psi_{2}\rangle$ For a given upper bound (termed error margin) on conditional error probabilities conditioned that, on measurement the state is declared as $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$, respectively, the probability of correct detection (i.e., the probability of declaring the state correctly) is maximized.
[17] Different 2 Unitary operation This work is a generalization of [18] for unitary operations with uniform prior probability. Closed-form expressions of maximum success probability for a given error margin are obtained for an irreducible group of unitary operations.
[14] More general $n$ Mixed states An extra inconclusive outcome is added while discriminating a set of $n$-mixed states. The goal is to maximize relative success probability (=1=1-postselected symmetric error) for the measurements for which the probability of inconclusive outcomes (=1=1-expected acceptance) is fixed.
[33] More general 2 Mixed states Minimum postselected error is obtained for states with known prior probability. An example measurement is given for which the minimum error is achieved.
Our work More general 2 Mixed stated Our work studies PSD of two mixed states with equal support, resulting in a non zero minimum postselected error as given in [33]. The set of all measurements is obtained for which the minimum postselected error is achieved, along with a parameterization of such measurements for complete characterization. If the support of the two states is not equal, it reduces to the USD case. Acceptance is obtained for each error-minimizing measurement along with the derivation of maximum achievable acceptance over the set of all error-minimizing measurements. If the support of the two states is not equal, it reduces to the USD case which is also studied as special case of PSD.
TABLE I: The system model and key contributions of relevant references [9, 13, 37, 4, 16, 42, 43, 7, 26, 28, 18, 17, 14, 33] in context of our work.
As mentioned above, USD is not always feasible. The work in [13, Theorem 2] showed that USD is feasible only when the two states in the pair of states have unequal support. Note that the support $\Pi_{\nu}$ of a state $\nu$ refers to the projection operator onto the eigenspace spanned by the set of eigenvectors corresponding to all nonzero eigenvalues of $\nu$. It is possible to generalize USD further to include the cases when the supports of the pair of states are equal. One such generalization is postselected discrimination (PSD), which utilizes measurements that can give one of the two states or reject as its outcome. Here, the reject outcome corresponds to not selecting any of the hypotheses (see Fig. 1 (c)) and declaring a failure of the test [33]. In other words, the decision is only made when the third outcome does not occur, which can be collectively termed selection of the test. Hence, such discrimination is termed postselected discrimination. Note that in contrast to USD, in PSD, a decision can be either incorrect or correct, resulting in a certain error post selection.
Figure 1: In all three subfigures, the black boxes on the left represent an unknown object in the state $\rho$ or $\sigma$. The colored boxes on the right represent the possible outcomes, where the green box corresponds to the accepted outcome, and the red box corresponds to the rejected outcome. The green line represents correct detection, the blue line shows incorrect detection, and the red line shows rejecting the test. Note that, in conventional hypothesis testing, only accepted outcomes were considered. Furthermore, from (b), it is clear that outcomes corresponding to blue lines are not allowed in USD.
The postselected probability of any event is defined as the probability of the event conditioned on the observation of one of the selected outcomes. Hence, in the symmetric PSD, the postselected symmetric error (See (4)) is defined as the probability of getting an incorrect outcome (corresponding to blue lines in Fig. 1 (c)) given that the test is selected (corresponding to both blue and green lines, which merge to green boxes, in Fig. 1 (c)). Recall that the event that the test is selected corresponds to the case when a decision is made and includes getting a correct or incorrect outcome. The minimum achievable value of postselected symmetric error is given in [33], along with an example measurement that achieves this value. However, a general characterization of all measurements that can achieve this error is not given in the prior work.
It is reasonable to wonder if, in postselected hypothesis testing, postselected symmetric error is the only performance metric and what constitutes an optimal measurement. To understand this, let us consider an example where two distinct measurements have the same postselected symmetric error, but different rejection probabilities. Here, rejection probability refers to the probability of obtaining the reject outcome, i.e., the probability of declaring the test failure and not making a decision. Essentially, even though the error is the same, one should prefer the one with a lower rejection probability. Further, a measurement that achieves the minimum value of error may have a high rejection probability. Such measurements are practically unusable. This possibility raises the question of how we can find measurements that minimize the probability of rejection and the postselected error. To address this concern, in our preliminary work [15], we define another performance metric, acceptance, as the probability that one of the first two outcomes occurs and a decision is made to select one of the two hypotheses. Hence, the acceptance for any state is given by the probability of observing an accepted outcome, given that the quantum object is in that state. In particular, acceptance for state $\rho$ (or $\sigma$) is the probability of events corresponding to green and blue lines originating from states $\rho$ (or $\sigma$, depending on the state under consideration), respectively, in Fig. 1 (c). It is intuitive to find a measurement that maximizes acceptance while minimizing the postselected error, which is the primary focus of the work.
Contributions: In this work, we consider symmetric postselected discrimination (PSD) for a pair of mixed states, $\rho$ and $\sigma$. We first consider the case where the two states have equal support, making USD infeasible. Hence, the PSD will result in a non-zero postselected error. We first determine the set of all measurements that minimize the postselected symmetric error. In particular, we represent error-minimizing measurements in a parametric form, allowing for a complete characterization of these measurements. Such parameterization represents a novel contribution compared to past literature, as it enables us to further refine the set of error-minimizing measurements to optimize any other performance metrics of interest. In particular, compute the maximum possible value of acceptance and provide an example for a measurement that achieves the maximum acceptance. This part of the work is entirely different than USD and is unrelated to the prior literature.
Next, we study USD as a special case of PSD where the supports of the two states are not equal. As discussed before, in this case, the postselected error vanishes, and therefore, the probability of a correct decision (also known as the success probability) is simply given as the expected acceptance. We first present the set of all error-minimizing (zero-error here) measurements along with their parameterized form, which satisfy the conditions given for USD measurements in [9, 13, 37, 4, 26, 28]. While most of these prior works have focused on maximizing success probability or obtaining general bounds on it (see [9, 13, 37, 4, 16, 26]), we provide closed-form expressions of the maximum acceptance for each state separately and show achievability by presenting one example of a measurement that achieves this maximum value of acceptance. The key contributions of our paper can be summarized as follows.
@@
    Characterization of the set of error-minimizing measurements - The main contribution of our work lies in determining the set of all error-minimizing measurements (i.e., measurements that can achieve the minimum value of the postselected symmetric error) in PSD and providing a parameterization to express these measurements for their complete characterization, which was not available in previous works. While the prior work [33] has presented the minimum value of postselected symmetric error in PSD, along with an example of such error-minimizing measurement, it does not provide a list of all error-minimizing measurements. It is not trivial to obtain this list, as [33] utilized the Lagrange duality approach to derive the minimum error, which does not provide any pointers towards finding the set of all optimal points. Finding the exact constraint on such measurement operators was a challenge, which our work resolved. Characterizing the complete set is essential to present any discussion on the optimality of measurements. To characterize error-minimizing measurements, we begin with noticing that any measurement that achieves the supremum given in (7) minimizes the error. We focus on finding the necessary and sufficient conditions for a measurement to achieve this supremum and thus obtain the complete set of measurements achieving the minimum error, as shown in the proof of Theorem 1, which is given in Appendix A. While deriving, we also obtain an alternative and novel derivation of the minimum error given in [33], which we have included in the same theorem which further validates our results.
@@
    Construction: unique and appropriate parameterization - We provide a construction of error-minimizing measurements based on parameters that can be varied appropriately to obtain all such measurements. A measurement constructed in this way is always an error-minimizing measurement. Furthermore, the parameterization is unique for each error-minimizing measurement in the sense that, changing the value of any of the parameters gives a different measurement. One key benefit of having such characterization is that we can refine these measurements to optimize other performance metrics, such as acceptance. Acceptance is a key metric proposed and maximized in this work for error-minimizing measurements. The parameterization provided us with a closed-form expression of acceptance, thus offering a direct approach to finding the maximum value of acceptance. The presented parameterization enables us to design an error-minimizing measurement to achieve any desired value of acceptance for each state. This observation is added as Remark 6 and Remark 14. The parameterization also helped us in finding the maximum acceptance easily, which is evident from the closed-form expression of the maximum acceptance in many cases.
@@
    A novel technique: subspace transformation for operators - In Theorem 1, we obtained the constraint on an arbitrary measurement given by a positive operator valued measure (POVM) $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ for it to be error-minimizing. We get these constraints on $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$, and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ as $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\Pi_{1})\text{ and }\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\Pi_{2}),$ where $\Pi_{1}$ and $\Pi_{2}$, obtained in Theorem 1, correspond to the projection operator on a pair of orthogonal subspaces. $\mathcal{P}(\Pi)$ denotes the set of all positive semidefinite operators having eigenvectors (corresponding to non-zero eigenvalues) lying in the subspace corresponding to the projection operator $\Pi$. At this stage, if we proceed in a conventional sense, we would have written an optimization problem as maximizing acceptance over the set of measurements satisfying these constraints. With that approach, unique parameterization would have been quite complicated. To simplify it further, we came up with a novel subspace-transformation that simplifies the constraints. We first observe that with these constraints on the operator, with some transformation, we can obtain another projection operator $\Pi_{3}$ and $\Pi_{4}$ such that the condition $\Lambda_{\rho}\in\mathcal{P}(\Pi_{3})$, and $\Lambda_{\sigma}\in\mathcal{P}(\Pi_{4})$ gives equivalent constraints. The equivalence is shown in Appendix B. To the best of our knowledge, this transformation, proved in Theorem B.1 and Theorem B.2 of Appendix B, has not been shown and used previously in this context. Using the proposed transformation and the constraints given in Theorem 1, we directly obtain the set of all error-minimizing measurements, which is given in Theorem 2. Therefore, we have obtained a parameterization of $\Lambda_{\rho}$, and $\Lambda_{\sigma}$ from Theorem 2.
@@
    Observations: properties of error minimizing measurements - After finding constraints on error minimizing measurements, we present the properties of error minimizing measurements in Corollary 1 and Corollary 2. One key observation provided is that, for the case with $\rho$ and $\sigma$ having equal support (i.e., $\Pi_{\rho}=\Pi_{\sigma}$), each any error minimizing measurement has one of the following three properties depending on the value of prior probabilities $p_{\rho}$ and $p_{\sigma}$.
    1. (i)
      If $p_{\rho}>p_{\rho}^{*}$ (i.e., $p_{\sigma}<p_{\sigma}^{*}$): Error-minimizing measurement never detects $\sigma$. It either detects $\rho$ or rejects the test.
    2. (ii)
      If $p_{\rho}<p_{\rho}^{*}$ (i.e., $p_{\sigma}>p_{\sigma}^{*}$): Error-minimizing measurement never detects $\rho$. It either detects $\sigma$ or rejects the test.
    3. (iii)
      If $p_{\rho}=p_{\rho}^{*}$ (i.e., $p_{\sigma}=p_{\sigma}^{*}$): There are error-minimizing measurements detecting either/both the hypotheses.
    Here, $p_{\rho}^{*}$ and $p_{\sigma}^{*}$ are certain constants given by (16). An exposition is provided in Fig 4.
@@
    PSD as a generalization of USD - While USD is feasible only if the two subspaces where each state lies are not equal (see Theorem 2 in [13]), PSD is a more general framework allowing discrimination with non-zero error. We also study USD as a special case of PSD for the case when the supports of the two states are not equal. We first present the set of all error-minimizing (which becomes zero-error in USD) measurements along with their parameterized form, which satisfy the conditions given for USD measurements in [9, 13, 37, 4, 26].
The notation needed across the paper is given below.